# StarEmbed: Benchmarking Time Series Foundation Models on Astronomical Observations of Variable Stars

### v1

@@ Abstract
[…]
1 Introduction
The adoption of time-series foundation models (TSFMs), with pretraining corpora that span commerce, finance, electricity, and traffic data, is proliferating due to their highly capable, general-purpose representation learning of time-variable signals (H. Zhou et al., 2021; Y. Nie et al., 2022; X. Yang et al., 2024; G. Woo et al., 2024). TSFMs are not trained on astronomical observations, however, and this omission is consequential because astronomical time series (“light curves”) present regimes that are rare in standard benchmarks: multiple variates, irregular time sampling, missing data, and heteroscedasticity (Figure 1). More specifically, there are frequent gaps of variable intervals in the observations (see Figure 1), and the presence of clouds, which change day-to-day and hour-by-hour, yields heteroskedastic uncertainties for the individual observations. At the same time, modern surveys such as the Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019) and the forthcoming Vera C. Rubin Observatory (Ž. Ivezić et al., 2019) generate peta-scale volumes of multi-band light curves, creating both a pressing need and a unique opportunity to evaluate TSFM generalization on real scientific data.
Stars that exhibit brightness variations over regular, periodic intervals (periodic variable stars) are astrophysically valuable as they are unique probes of stellar interiors and evolution, galactic structure, and can be used to measure the distance to nearby galaxies (e.g., M. W. Feast & A. R. Walker, 1987; G. Clementini et al., 2003; K. Genovali et al., 2014; M. Catelan & H. A. Smith, 2015; V. Ripepi et al., 2017). Dozens of types of periodic variable stars exist, and modern astronomical surveys, like ZTF, have produced an avalanche of light curves ($\sim$109 stars each with $\sim$103 observations over 7 yr from ZTF alone). These light curves are multi-variate because observations are conducted with a filter (or “passband”) placed along the focal path of the telescope, limiting the image to only light from a specific wavelength range. Thus, light curves contain both brightness and “color” information (i.e., the relative brightnesses across passbands), enabling inference of physical properties of the source. The abundance of astronomical time series data will dramatically accelerate as the recently-commissioned Vera C. Rubin Observatory (Ž. Ivezić et al., 2019) will discover >>108 variable stars while monitoring >>$10$ stars over the course of a decade. Despite this abundance of astronomical light curves, there is no standardized benchmark for assessing time-series embeddings in this domain. The absence of common datasets, class sets, and train-test splits has hindered fair, reproducible comparisons and obscured whether domain-specific pipelines outperform generic representations from foundation models (cf., R. Pan et al., 2024).
@@ 1 Introduction
The remainder of this paper describes related works and the models we benchmark (Section 2); introduces the ZTF data set (Section 3), provides our benchmark methodology (Section 4), and presents the benchmark results (Section 5) before discussing our concluding thoughts (Section 6).
@@
2 Related Works and Models
@@
Major recent investments in time-domain astronomy have generated incredibly large datasets that naturally lend themselves to machine learning methods. The classification of periodic variable stars has been a problem of significant interest for centuries, as these sources provide direct insight into many facets of stellar astrophysics. As such, both pre- (e.g., J. Debosscher et al., 2007) and post-deep learning models (e.g., D. Moreno-Cartagena et al., 2025) have been applied to this problem. We summarize the embedding models and the baseline below. We aim to assess the zero-shot generalization capabilities of the pre-trained TSFMs, so we do not fine-tune them on our data.
2.1 Supervised Classifiers
@@
The first machine learning models to classify variable stars used manually engineered features combined with classical models such as support vector machines (J. Debosscher et al., 2007) or gradient boosted decision trees (K. Boone, 2019). J. W. Richards et al. (2011) achieved SOTA performance with 52 extracted features (including Fourier coefficients, variability amplitude, skewness, etc.) and a random forest (RF) classifier. Feature extraction varies from study to study, though some have attempted to standardize this step (I. Nun et al., 2015; K. L. Malanchev et al., 2021a). Recent work has introduced deep-learning methods to eliminate explicit feature engineering using a wide range of architectures including recurrent neural networks (RNNs) (D. Muthukrishna et al., 2019; I. Becker et al., 2020; V. G. Shah et al., 2025) and transformers (G. Cabrera-Vives et al., 2024; D. Moreno-Cartagena et al., 2025). These efforts, however, do not perform meaningfully better than the established hand-crafted feature extraction baseline: accuracies of RNN models are $\pm 1-3\%$ of hand-crafted features across multiple variable star datasets (see, e.g., B. Naul et al., 2018). As a result, we choose to use hand-crafted features to establish our baseline performance.
@@
Baseline Model: For this work, we first extract features using the FATS (I. Nun et al., 2015) and light_curve (K. L. Malanchev et al., 2021a) software packages. Example features include: the best-fit Lomb-Scargle (N. R. Lomb, 1976; J. D. Scargle, 1982) period, the scatter, the skewness, the kurtosis, and other metrics. In total, we define 69 features per passband, yielding a total embedding size of 138 for the two-passband ZTF data (see Appendix A for a full feature list with explanations). We normalize each feature to have zero mean and unit variance. While very effective, hand-crafted features rely heavily on domain knowledge, can be brittle to data quality issues, and are expensive to compute.
2.2 Astrophysics Embedding models
@@
With a high cost to obtain labels for astronomical sources, there has been a growing interest in using semi-supervised approaches to learn general representations of the data to later perform downstream tasks. Recent approaches include variational autoencoders (V. A. Villar et al., 2020), sparse autoencoders (S. Dillmann et al., 2025), and contrastive learning (G. Zhang et al., 2024), but they are typically limited to a single class (e.g., supernovae). A few foundation models for astronomy attempt to produce useful representations of light curves, such as FALCO (X. Zuo et al., 2025) and Astromer (C. Donoso-Oliva et al., 2023, 2025). Unlike FALCO, Astromer-1 and Astromer-2 are designed to apply to observations from any observatory (C. Donoso-Oliva et al., 2023, 2025), and thus, we adopt the Astromer models as an astronomy-specific foundation model.
@@
Astromer (C. Donoso-Oliva et al., 2023) is a transformer-based model to generate informative embedded representations of light curves. Astromer-1 was pre-trained using self-supervised learning on 1.5 million single-band light curves from the MACHO survey (C. Alcock et al., 2000). The model’s output is a fixed-length embedding, and, as recommended by the creators, we use the 256-dimensional embedding from the final attention layer produced using the publicly released weights. Astromer-1 was trained to reconstruct masked portions of the input sequence (i.e., masked time series modeling). Astromer-2 (C. Donoso-Oliva et al., 2025) increases the number of model parameters from 0.660.66M to 5.45.4M and adopts an uncertainty-weighted loss function for pretraining. The Astromer models represent the SOTA domain-specific model and serve as a prime candidate for testing whether astronomy-specific pre-training yields discernible benefits relative to general TSFMs.
2.3 Time series foundation models
@@
TSFMs have been shown to consistently outperform the traditional one-dataset-per-model schema in multiple fields, including finance, climate science, and commerce (e.g., Z. Yue et al., 2022; G. Woo et al., 2024; A. F. Ansari et al., 2024). Astronomy, despite having an enormous collection of light curves, has yet to examine the potential of TSFMs, which may prove transformative in our ability to accomplish multiple downstream tasks. Furthermore, large time-domain surveys provide a unique opportunity to evaluate TSFMs with minimal risk of data leakage because astronomical light curves are not included in any of the training corpora. This benchmark therefore provides a new test for how TSFMs transfer to an unseen domain.
@@
Moirai (G. Woo et al., 2024) is designed to be a single foundation model that can forecast virtually any time-series, regardless of sampling frequency, dimensionality, or distribution. It pairs a multi-patch-size projection scheme (i.e., handling minute- to year-scale data), an any-variate attention mechanism that scales to arbitrary numbers of variables, and a flexible mixture-distribution output head for calibrated probabilistic forecasts. Trained on LOTSA (G. Woo et al., 2024), an open archive of 27 billion observations spanning nine domains, Moirai’s Small/Base/Large variants deliver SOTA accuracy in both in-distribution and zero-shot settings, often outperforming models that are fully fine-tuned for a particular dataset.
@@
Chronos (A. F. Ansari et al., 2024) is another pre-trained time series model showing comparable or even better results than Moirai. It treats forecasting as a language-modeling problem: Chronos scales and quantizes real-valued time-series into a fixed vocabulary, then trains off-the-shelf Transformer language models (T5-style models with 20M to 710M parameters) with an ordinary cross-entropy loss. Augmented by TSMixup and Gaussian-process–generated synthetic data, Chronos is pre-trained on a large collection of public datasets and evaluated on 42 benchmarks. The resulting models deliver strong probabilistic forecasts—significantly ahead of classical and deep-learning baselines on in-domain data and in zero-shot settings, showing that “language of time-series” tokenization alone is enough to build a competitive universal forecaster.
2.4 Random Embeddings As a Sanity Check Baseline
@@
3 Dataset
The benchmark dataset includes multi-variate time-series observations of periodic variable stars. The flux is presented in magnitudes (an astronomy specific unit), while the time is recorded as the modified Julian date. The observations are from ZTF, which repeatedly scans all stars visible from the Northern hemisphere every few days. ZTF observes in three different passbands, $g$, $r$, and ii [^1] Most ZTF sources have very few or no observations in $i$ band and we therefore exclude it from our analysis. (see Figure˜1) roughly corresponding to visible green, visible red, and (outside the visible) infrared light, respectively (R. Dekany et al., 2020). We use observations from ZTF data release 23 (DR23) which spans a duration of $\sim$6.5 yr and contains billions of light curves.
@@ 3 Dataset
For training, we instead adopt the Catalina Surveys Periodic Variable Star Catalog (CSPVS; A. J. Drake et al., 2014a), a human-labeled catalog of periodic variables discovered by the Catalina Real-Time Transient Survey (A. J. Drake et al., 2009). We extract ZTF light curves for each CSPVS star. Stars that lack a ZTF light curve are omitted; we also remove (i) observations flagged as “bad” in ZTF DR23; (ii) light curves with <32<32 total observations; (iii) light curves that lack both $g$ and $r$ observations; and (iv) light curves from classes with fewer than 350 total examples. The resulting dataset contains $\sim$40,000 ZTF light curves of expert-labeled periodic variable stars across seven classes. See Appendix B for a detailed astrophysical description of each class.
@@
3.1 ZTF Dataset in Context: Relevance to Upcoming Observatories
Variable star science has an expansive scope that extends beyond ZTF and periodic variables; the StarEmbed benchmark is designed to allow for the addition of new datasets and metrics in future expansions. This flexibility is crucial to the long-term health of this benchmark as numerous new time-domain surveys, like LSST, will begin in the coming years. Each new survey has unique observational capabilities and priorities that will affect the resulting embeddings and downstream task performance. As the largest astronomical time-domain experiment to date, ZTF is an apt choice for building a preparatory benchmark dataset that has already been used to explore emerging areas like multi-modality (e.g., D. A. Duev & S. J. van der Walt, 2021; R. Carrasco-Davis et al., 2021; A. Gagliano et al., 2023; N. Rehemtulla et al., 2024) and transformers (e.g., T. Allam et al., 2023; G. Zhang et al., 2024).
All current and future datasets associated with the StarEmbed benchmark are public or will be made public. ZTF DR23 data can be accessed through the Caltech Infrared Processing and Analysis Center [^2] https://irsa.ipac.caltech.edu/Missions/ztf.html; the CSPVS catalog is available via VizieR [^3] https://cdsarc.cds.unistra.fr/viz-bin/cat/J/ApJS/213/9 (A. J. Drake et al., 2014b); and our selection of ZTF light curves for CSPVS stars will be available on Hugging Face. Our publicly available dataset includes the necessary metadata and a permissive license for reuse. No personal or sensitive information is present in these datasets (they consist of astronomical observations).

### v2

@@ Abstract
I Introduction
The adoption of time-series foundation models (TSFMs), with pretraining corpora that span commerce, finance, electricity, and traffic data, is proliferating due to their highly capable, general-purpose representation learning of time-variable signals (Zhou et al., 2021a; Nie et al., 2022; Yang et al., 2024; Woo et al., 2024). TSFMs are not trained on astronomical observations, however, and this omission is consequential because astronomical time series (“light curves”) present regimes that are rare in standard benchmarks: multiple variates, irregular time sampling, missing data, and heteroscedasticity (Figure 1). More specifically, there are frequent gaps of variable intervals in the observations (see Figure 1), and the presence of clouds, which change day-to-day and hour-by-hour, yields heteroskedastic uncertainties for the individual observations. At the same time, modern surveys such as the Zwicky Transient Facility (ZTF; Bellm et al., 2019) and the forthcoming Vera C. Rubin Observatory (Ivezić et al., 2019) generate peta-scale volumes of multi-band light curves, creating both a pressing need and a unique opportunity to evaluate TSFM generalization on real scientific data.
Stars that exhibit brightness variations over regular, periodic intervals (periodic variable stars) are astrophysically valuable as they are unique probes of stellar interiors and evolution, galactic structure, and can be used to measure the distance to nearby galaxies (e.g., Feast and Walker, 1987; Clementini et al., 2003; Genovali et al., 2014; Catelan and Smith, 2015; Ripepi et al., 2017). Dozens of types of periodic variable stars exist, and modern astronomical surveys, like ZTF, have produced an avalanche of light curves ($\sim$109 stars each with $\sim$103 observations over 7 yr from ZTF alone). These light curves are multi-variate because observations are conducted with a filter (or “passband”) placed along the focal path of the telescope, limiting the image to only light from a specific wavelength range. Thus, light curves contain both brightness and “color” information (i.e., the relative brightnesses across passbands), enabling inference of physical properties of the source. The abundance of astronomical time series data will dramatically accelerate as the recently-commissioned Vera C. Rubin Observatory (Ivezić et al., 2019) will discover >>108 variable stars while monitoring >>$10$ stars over the course of a decade. Despite this abundance of astronomical light curves, there is no standardized benchmark for assessing time-series embeddings in this domain. The absence of common datasets, class sets, and train-test splits has hindered fair, reproducible comparisons and obscured whether domain-specific pipelines outperform generic representations from foundation models (cf., Pan et al., 2024b).
@@ I Introduction
The remainder of this paper describes related works and the models we benchmark (Section II); introduces the ZTF data set (Section III), provides our benchmark methodology (Section IV), and presents the benchmark results (Section V) before discussing our concluding thoughts (Section VI).
@@
II Related Works and Models
@@
Major recent investments in time-domain astronomy have generated incredibly large datasets that naturally lend themselves to machine learning methods. The classification of periodic variable stars has been a problem of significant interest for centuries, as these sources provide direct insight into many facets of stellar astrophysics. As such, both pre- (e.g., Debosscher et al., 2007) and post-deep learning models (e.g., Moreno-Cartagena et al., 2025) have been applied to this problem. We summarize the embedding models and the baseline below. We aim to assess the zero-shot generalization capabilities of the pre-trained TSFMs, so we do not fine-tune them on our data.
II.1 Supervised Classifiers
@@
The first machine learning models to classify variable stars used manually engineered features combined with classical models such as support vector machines (Debosscher et al., 2007) or gradient boosted decision trees (Boone, 2019). Richards et al. (2011) achieved SOTA performance with 52 extracted features (including Fourier coefficients, variability amplitude, skewness, etc.) and a random forest (RF) classifier. Later work involved using goodness-of-fit metrics from fitting template physical models to the time series as additional inputs into a tree-based classifier Sesar et al. (2017). Feature extraction varies from study to study, though some have attempted to standardize this step (Nun et al., 2015; Kim and Bailer-Jones, 2016; Malanchev et al., 2021a). Recent work has introduced deep-learning methods to eliminate explicit feature engineering using a wide range of architectures including recurrent neural networks (RNNs) (Muthukrishna et al., 2019; Becker et al., 2020; Shah et al., 2025) and transformers (Cabrera-Vives et al., 2024; Moreno-Cartagena et al., 2025). These efforts, however, do not perform meaningfully better than the established hand-crafted feature extraction baseline: accuracies of RNN models are $\pm 1-3\%$ of hand-crafted features across multiple variable star datasets (see, e.g., Naul et al., 2018). As a result, we choose to use hand-crafted features to establish our baseline performance.
@@
Baseline Model: For this work, we first extract features using the FATS (Nun et al., 2015) and light_curve (Malanchev et al., 2021a) software packages. Example features include: the best-fit Lomb-Scargle (Lomb, 1976; Scargle, 1982) period, the scatter, the skewness, the kurtosis, and other metrics. In total, we define 69 features per passband, yielding a total embedding size of 138 for the two-passband ZTF data (see Appendix A for a full feature list with explanations). We normalize each feature to have zero mean and unit variance. While very effective, hand-crafted features rely heavily on domain knowledge, can be brittle to data quality issues, and are expensive to compute.
II.2 Astrophysics Embedding models
@@
With a high cost to obtain labels for astronomical sources, there has been a growing interest in using semi-supervised approaches to learn general representations of the data to later perform downstream tasks. Recent approaches include variational autoencoders (Villar et al., 2020), sparse autoencoders (Dillmann et al., 2025), and contrastive learning (Zhang et al., 2024), but they are typically limited to a single class (e.g., supernovae). A few foundation models for astronomy attempt to produce useful representations of light curves, such as FALCO (Zuo et al., 2025) and Astromer (Donoso-Oliva et al., 2023, 2025). Unlike FALCO, Astromer-1 and Astromer-2 are designed to apply to observations from any observatory (Donoso-Oliva et al., 2023, 2025), and thus, we adopt the Astromer models as an astronomy-specific foundation model.
@@
Astromer (Donoso-Oliva et al., 2023) is a transformer-based model to generate informative embedded representations of light curves. Astromer-1 was pre-trained using self-supervised learning on 1.5 million single-band light curves from the MACHO survey (Alcock et al., 2000). The model’s output is a fixed-length embedding, and, as recommended by the creators, we use the 256-dimensional embedding from the final attention layer produced using the publicly released weights. Astromer-1 was trained to reconstruct masked portions of the input sequence (i.e., masked time series modeling). Astromer-2 (Donoso-Oliva et al., 2025) increases the number of model parameters from 0.660.66M to 5.45.4M and adopts an uncertainty-weighted loss function for pretraining. The Astromer models represent the SOTA domain-specific model and serve as a prime candidate for testing whether astronomy-specific pre-training yields discernible benefits relative to general TSFMs.
II.3 Time series foundation models
@@
TSFMs have been shown to consistently outperform the traditional one-dataset-per-model schema in multiple fields, including finance, climate science, and commerce (e.g., Yue et al., 2022; Woo et al., 2024; Ansari et al., 2024). With strong performance that scales with model and data set size, they are a promising tool for driving the future of AI for time series Edwards et al. (2024); Pan et al. (2024a). Astronomy, despite having an enormous collection of light curves, has yet to examine the potential of TSFMs, which may prove transformative in our ability to accomplish multiple downstream tasks. Furthermore, large time-domain surveys provide a unique opportunity to evaluate TSFMs with minimal risk of data leakage because astronomical light curves are not included in any of the training corpora. This benchmark therefore provides a new test for how TSFMs transfer to an unseen domain.
Moirai (Woo et al., 2024)
@@
is designed to be a single foundation model that can forecast virtually any time-series, regardless of sampling frequency, dimensionality, or distribution. It pairs a multi-patch-size projection scheme (i.e., handling minute- to year-scale data), an any-variate attention mechanism that scales to arbitrary numbers of variables, and a flexible mixture-distribution output head for calibrated probabilistic forecasts. Trained on LOTSA (Woo et al., 2024), an open archive of 27 billion observations spanning nine domains, Moirai’s Small/Base/Large variants deliver SOTA accuracy in both in-distribution and zero-shot settings, often outperforming models that are fully fine-tuned for a particular dataset.
@@
Chronos (Ansari et al., 2024) is another pre-trained time series model showing comparable or even better results than Moirai. It treats forecasting as a language-modeling problem: Chronos scales and quantizes real-valued time-series into a fixed vocabulary, then trains off-the-shelf Transformer language models (T5-style models with 20M to 710M parameters) with an ordinary cross-entropy loss. Augmented by TSMixup and Gaussian-process–generated synthetic data, Chronos is pre-trained on a large collection of public datasets and evaluated on 42 benchmarks. The resulting models deliver strong probabilistic forecasts—significantly ahead of classical and deep-learning baselines on in-domain data and in zero-shot settings, showing that “language of time-series” tokenization alone is enough to build a competitive universal forecaster.
II.4 Random Embeddings As a Sanity Check Baseline
@@
III Dataset
The benchmark dataset includes multi-variate time-series observations of periodic variable stars. The flux is presented in magnitudes (an astronomy specific unit), while the time is recorded as the modified Julian date.
Figure 2: Statistical analysis of our train and test split data. Left: histogram of sampling intervals (days between consecutive measurements). Right: histogram of observational uncertainties. Sampling frequency varies across >>4 orders of magnitude, reflecting the significant challenge of irregular sampling in our light curves. Uncertainties on the target also vary across multiple orders of magnitude. These features are critical properties of light curves which are rarely to never seen in TSFM pretraining data sets.
The observations are from ZTF, which repeatedly scans all stars visible from the Northern hemisphere every few days. ZTF observes in three different passbands, $g$, $r$, and ii [^1] Most ZTF sources have very few or no observations in $i$ band and we therefore exclude it from our analysis. (see Figure˜1) roughly corresponding to visible green, visible red, and (outside the visible) infrared light, respectively (Dekany et al., 2020). We use observations from ZTF data release 23 (DR23) which spans a duration of $\sim$6.5 yr and contains billions of light curves. Figure 2 shows a statistical analysis of this data set, highlighting the challenges inherent to these astronomical time series. The left panel shows the sampling frequency or the distribution of time elapsed between two measurements of a given star. This demonstrates the significant irregularity in the temporal sampling of this data. The sparsity at high frequency sampling reflects the limited scientific utility of very high frequency observing, and the gap at frequencies of two per day reflects the fact that observing from the ground can only occur at nighttime and it is always daytime 0.5 days after an observation is taken. The sampling frequency is very consistent across the $g$ and $r$ bands and across the train and test splits. The right panel shows the distribution of measured magnitude uncertainties. Many factors can dramatically affect these uncertainties, and this is reflected in their large spread. We observe that the $g$ and $r$ band uncertainty distributions are slightly different; $r$ band uncertainties are more centrally concentrated and $g$ band uncertainties have greater extremes. We also observe a slight over-density of $g$ band measurements with very small uncertainties in the train split. Noting the logarithmic scale on the y-axis, this is not a concern as these represent only a very small number of observations.
We also note that these features are in stark contrast to those in time series data sets commonly used for TSFM pretraining. LOTSA Woo et al. (2024), e.g., only includes time series without spread in their sampling frequencies, i.e. only regularly sampled time series, and nearly all time series in LOTSA do not include uncertainties on the target or multiple variates. Therefore, this astronomical time series data represents a new, special out-of-domain test for the TSFMs we explore here.
@@ III Dataset
For training, we instead adopt the Catalina Surveys Periodic Variable Star Catalog (CSPVS; Drake et al., 2014a), a human-labeled catalog of periodic variables discovered by the Catalina Real-Time Transient Survey (Drake et al., 2009). We extract ZTF light curves for each CSPVS star. Stars that lack a ZTF light curve are omitted; we also remove (i) observations flagged as “bad” in ZTF DR23; (ii) light curves with <32<32 total observations; (iii) light curves that lack both $g$ and $r$ observations; and (iv) light curves from classes with fewer than 350 total examples. The resulting dataset contains $\sim$40,000 ZTF light curves of expert-labeled periodic variable stars across seven classes. See Appendix B for a detailed astrophysical description of each class.
@@
III.1 ZTF Dataset in Context: Relevance to Upcoming Observatories
Variable star science has an expansive scope that extends beyond ZTF and periodic variables; the StarEmbed benchmark is designed to allow for the addition of new datasets and metrics in future expansions. This flexibility is crucial to the long-term health of this benchmark as numerous new time-domain surveys, like LSST, will begin in the coming years. Each new survey has unique observational capabilities and priorities that will affect the resulting embeddings and downstream task performance. As the largest astronomical time-domain experiment to date, ZTF is an apt choice for building a preparatory benchmark dataset that has already been used to explore emerging areas like multi-modality (e.g., Duev and van der Walt, 2021; Carrasco-Davis et al., 2021; Gagliano et al., 2023; Rehemtulla et al., 2024) and transformers (e.g., Allam et al., 2023; Zhang et al., 2024).
All current and future datasets associated with the StarEmbed benchmark are public or will be made public. ZTF DR23 data can be accessed through the Caltech Infrared Processing and Analysis Center [^2] https://irsa.ipac.caltech.edu/Missions/ztf.html; the CSPVS catalog is available via VizieR [^3] https://cdsarc.cds.unistra.fr/viz-bin/cat/J/ApJS/213/9 (Drake et al., 2014b); and our selection of ZTF light curves for CSPVS stars will be available on Hugging Face. Our publicly available dataset includes the necessary metadata and a permissive license for reuse. No personal or sensitive information is present in these datasets (they consist of astronomical observations).